OpenStudy (anonymous):

ok my problem is 6+2(x-7)=10-3(x-3) 6+2x-14=10-3x+9 -8+2x=19-3x -27=-5x and the answer is x=27/5 but what my new problem is that i need to do a check step and i have no clue

OpenStudy (stormfire1):

Plug in 27/5 for x in the original equation and see if both sides are still equivalent.

OpenStudy (anonymous):

i put in 27/5 and got 5.4

OpenStudy (stormfire1):

Ok, so you got 5.4 which is correct. Now plug that into the original equation and see if both sides are equivalent\[6+2(5.4-7)=10-3(5.4-3)\]

OpenStudy (anonymous):

yea i just divide 27 by 5 i am sorry i have no clue what to do iam dumb

OpenStudy (anonymous):

i try to do alot on my own but if i fail this class then its 1000 out of my own pocket

OpenStudy (stormfire1):

Not dumb...just learning. Your problem asked you to solve for x and you found it to be 5.4. If 5.4 is the correct answer then both sides of the equation will come to the same value (ie: they are equivalent) and your answer has been shown to be correct. \[6+2(5.4-7)=2.8\]\[10-3(5.4-3)=2.8\]

OpenStudy (stormfire1):

Since both answers are 2.8, your answer of 27/5 (or 5.4) is valid.

OpenStudy (anonymous):

ok that sounds like i get that some

OpenStudy (stormfire1):

If you had come up with a different value of x then most likely the above tests would have shown two different answers...telling you that you did something wrong. This is all just a way to check your answer.

OpenStudy (anonymous):

ohhh but why is it inportant taht we do that

OpenStudy (stormfire1):

Well, if you didn't do that how would you know that you performed the algebra correctly?

OpenStudy (stormfire1):

For simple problems you can probably just double-check your algebra and be fine...but when the problems get a lot harder, it's nice to be able to check your answers. Imagine what could happen if you built a rocket based on unchecked math :P

OpenStudy (anonymous):

ok i can understand why when you explan it like that